OpenStudy (anonymous):

Need help! -5= 8m+12 ------- 4 Thanks!

OpenStudy (anonymous):

\[\frac{ 8m+12 }{ 4 }= -5\]

OpenStudy (anonymous):

cross multiply subtract both side by 12 divide both side by 8 and that is how u get your answer

OpenStudy (anonymous):

@Daniel56k How do I cross cancel this?

OpenStudy (unklerhaukus):

multiply both sides by 4

OpenStudy (anonymous):

-5 multiply by 4 is equal to 8m+12

OpenStudy (anonymous):

Ok so 8m+48 now?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

no

OpenStudy (unklerhaukus):

\[\frac{ 8m+12 }{ 4 }= -5\\ \frac{ 8m+12 }{ 4 }\times4= -5\times4\\ \frac{ 8m+12 }{ \cancel{4} }\cancel{\times4}= -5\times4\\ \qquad\dots\qquad=\qquad\dots\]

OpenStudy (anonymous):